Find the height 'h' such as the following 2D force system is in equilibrium:

W = 450 N

d = 800 mm

p=1.06W
Answer #1

or - W = 450N d = 800 mm w p=106W. p= 477 N. By equilibrium of up. forces, we get; Tsino =P - (in hos and, T=W - (ii) From (i